# https://mvsgaussian.github.io/

### v1

@@ 2 Related Work
Generalizable NeRF. By implicitly representing scenes as continuous color and density fields using MLPs, Neural Radiance Fields (NeRF) achieve impressive rendering results with volume rendering techniques. Follow-up works [5, 42, 2, 27, 41, 35] extend it to various tasks and achieve promising results. However, they all require time-consuming per-scene optimization. To address this issue, some generalizable NeRF methods have been proposed. The general paradigm involves encoding features for each 3D point and then decoding these features to obtain volume density and radiance. According to the encoded features, generalizable NeRFs can be categorized into appearance features [50], aggregated multi-view features [36, 19, 21, 33], cost volume-based features [6, 23, 19, 21], and correspondence matching features [8]. Despite achieving significant results, performance remains limited, with slow optimization and rendering speeds.
3D Gaussian Splatting (3D-GS) utilizes anisotropic Gaussians to explicitly represent scenes and achieves real-time rendering through differentiable rasterization. Motivated by this, several studies have applied it to various tasks, e.g. editing [7, 3], dynamic scenes [46, 24, 40], and avatars [15, 14, 29]. However, the essence of Gaussian Splatting still lies in overfitting the scene. To remedy this, a few concurrent works make initial attempts to generalize Gaussian Splatting to unseen scenes. The goal of Generalizable Gaussian Splatting is to predict Gaussian parameters in a feed-forward manner instead of per-scene optimization. PixelSplat [4] addresses scale ambiguity by leveraging an epipolar Transformer to encode features and subsequently decode them into Gaussian parameters. However, it focuses on image pairs as input and the Transformer incurs significant computational costs. GPS-Gaussian [52] draws inspiration from stereo matching and performs epipolar rectification and disparity estimation on input image pairs. However, it focuses on human novel view synthesis and requires ground-truth depth maps. Spatter Image [32] introduces a single-view 3D reconstruction approach. However, it focuses on object-centric reconstruction rather than generalizing to unseen scenes. Overall, these methods are constrained by inefficiency, limited to object reconstruction, and restricted to either image pairs or a single view. To this end, in this paper, we aim to study an efficient generalizable Gaussian Splatting for novel view synthesis in unseen general scenes.
@@ 3 Preliminary
$\small C=\sum_{i}c_{i}\alpha_{i}{\prod_{j=1}^{i-1}(1-\alpha_{i})}\,,$ (2)
where $c_{i}$ is the color of each point, defined by spherical harmonics (SH) coefficients. The density $\alpha_{i}$ is computed by the multiplication of 2D Gaussians and a learnable point-wise opacity.
During optimization, the learnable attributes of each Gaussian are updated through gradient descent, including 1) a 3D position $\mu\in\mathbb{R}^{3}$, 2) a scaling vector $s\in\mathbb{R}_{+}^{3}$, 3) a quaternion rotation vector $r\in\mathbb{R}^{4}$, 4) a color defined by SH $c\in\mathbb{R}^{k}$ (where k is the freedom), and 5) an opacity $\alpha\in[0,1]$. Additionally, an adaptive density control module is introduced to improve rendering quality, comprising mainly the following three operations: 1) split into smaller Gaussians if the magnitude of the scaling exceeds a threshold, 2) clone if the magnitude of the scaling is smaller than a threshold, and 3) prune Gaussians with excessively small opacity or overly large scaling magnitudes.
@@ 4 MVSGaussian
Figure 2: The overview of generalizable Gaussian Splatting framework. We first extract features $\{f_{i}\}_{i=1}^{N}$ from input source views $\{I_{i}\}_{i=1}^{N}$ using FPN. These multi-view features are then aggregated into a cost volume, regularized by 3D CNNs to produce depth estimations. Subsequently, for each 3D point at the estimated depth, we utilize a pooling network to aggregate warped source features, obtaining the aggregated feature $f_{v}$. This feature is then enhanced using a 2D UNet, yielding the enhanced feature $f_{g}$. $f_{g}$ is decoded into Gaussian parameters for splatting, while $f_{v}$ is decoded into volume density and radiance for depth-aware volume rendering with a single sampled point. Finally, the two rendered images are averaged to produce the final rendered result.
@@ 4.3 Consistent Aggregation for Per-Scene Optimization
Figure 3: Consistent aggregation. With the depth maps and Gaussian point clouds produced by the generalizable model, we first conduct multi-view geometric consistency checks on the depth maps to derive masks for filtering out unreliable points. Subsequently, the filtered point clouds are concatenated to construct a point cloud, serving as the initialization for per-scene optimization.
@@ 5.2 Generalization Results
5.3 Per-Scene Optimization Results
The quantitative results after per-scene optimization are reported in Table 3. For per-scene optimization, one strategy is to optimize the entire pipeline, similar to NeRF-based methods. Another approach is to optimize only the initial Gaussian point cloud provided by the generalizable model. When optimizing the entire pipeline, Our method can achieve better performance with faster inference speeds compared to previous generalizable NeRF methods. it demonstrates comparable results to NeRF, showcasing the robust representation capabilities of our method. In contrast, optimizing only the Gaussians can significantly improve optimization and rendering speed because it eliminates the time-consuming feed-forward neural network. Moreover, performance can benefit from the adaptive density control module described in Sec. 3. Due to the excellent initialization provided by the generalizable model and the effective aggregation strategy, we achieve optimal performance within a short optimization period, approximately one-tenth of that of 3D-GS [16]. Especially on the Real Forward-facing dataset, our method achieves superior performance with only $45$ seconds of optimization, compared to $10$ minutes for 3D-GS and $10$ hours for NeRF [26]. Additionally, our method’s inference speed is comparable to that of 3D-GS and significantly outperforms NeRF-based methods. As shown in Fig. 5, our method is capable of producing high-fidelity views with finer details.
5.4 Ablations and Analysis
@@
Table 5: Comparison of different aggregation strategies. We report the quantitative results obtained with different strategies on the Real Forward-facing dataset [25]. For downsampling aggregation, we employ widely-used voxel downsampling with a voxel size set to $2$. The iteration number for all aggregation strategies is set to $2$k.
@@ 6 Conclusion
In this paper, we present MVSGaussian, a novel generalizable Gaussian Splatting approach for reconstructing scene representations from MVS. Specifically, we leverage MVS to encode geometry-aware features, establishing a pixel-aligned Gaussian representation. Additionally, we propose a hybrid Gaussian rendering approach that integrates efficient depth-aware volume rendering to enhance generalization. Apart from its remarkable generalizability, our model can be easily fine-tuned for specific scenes. To facilitate fast optimization, we introduce a multi-view geometric consistent aggregation strategy to provide high-quality initialization. Compared with generalizable NeRFs, which typically require minutes of fine-tuning and seconds of rendering per image, MVSGaussian achieves real-time rendering with superior synthesis quality. Moreover, compared with the vanilla 3D-GS, MVSGaussian achieves better view synthesis with reduced training computational cost. Extensive experiments validate that MVSGaussian achieves state-of-the-art performance with convincing generalizability, real-time rendering speed, and fast per-scene optimization.
@@ Appendix 0.A Implementation and Network Details
Implementation Details. Following ENeRF [19], we partition the DTU [1] dataset into $88$ training scenes and $16$ test scenes. We train the generalizable model on four RTX $3090$ GPUs using the Adam [17] optimizer, with an initial learning rate set to $45e-$ 4. The learning rate is halved every $50$k iterations. During the training process, we select $2$, $3$, and $4$ source views as inputs with respective probabilities of $0.1$, $0.8$, and $0.1$. For evaluation, we follow the criteria established in prior works such as ENeRF [19] and MVSNeRF [6]. Specifically, for the DTU test set, segmentation masks are employed to evaluate performance, defined based on the availability of ground-truth depth at each pixel. For Real Forward-facing dataset [25], where the marginal region of images is typically invisible to input images, we evaluate the $80\%$ area in the center of the images. This evaluation methodology is also applied to the Tanks and Temples dataset [18]. The image resolutions of the DTU, the Real Forward-facing, the NeRF Synthetic [26], and the Tanks and Temples datasets are $512\times 640$, $640\times 960$, $800\times 800$, and $640\times 960$ respectively. As discussed in Sec. 4.3 of the main text, we employ a consistency check to filter out noisy points for high-quality initialization. Specifically, we apply a dynamic consistency checking algorithm [43, 22], the details of which are provided in Algorithm 1. The predefined thresholds $\{\theta_{p}(n)\}_{n=1}^{N_{\theta}}$ are set to $\{\frac{n}{4}\}_{n=1}^{N_{\theta}}$, and $\{\theta_{d}(n)\}_{n=1}^{N_{\theta}}$ are set to $\{\frac{n}{10}\}_{n=1}^{N_{\theta}}$. For 3D-GS [16], following [53], we use COLMAP [30] to reconstruct the point cloud from the working set (training views) as initialization. Specifically, we employ COLMAP’s automatic reconstruction to achieve the reconstruction of sparse point clouds. Some examples are shown in Fig. 7.
@@ 21       $Mask\leftarrow Mask\cup Mask_{n}$
Appendix 0.B Additional Ablation Experiments
Inference Speed Analysis. As shown in Table 1 of the main text, the inference speed of our generalizable model is $21.5$FPS. Here, we present the inference time breakdown result in Table 10. The primary time overhead comes from the neural network, while the subsequent rendering process incurs minimal time overhead. Therefore, we discard the neural network component during the per-scene optimization stage, resulting in a significant increase in speed.
Appendix 0.C More Qualitative Results
@@
Appendix 0.D Per-scene Breakdown
As shown in Tables 111413, and 12, we present the per-scene breakdown results of DTU [1], NeRF Synthetic [26], Real Forward-facing [25], and Tanks and Temples [18] datasets. These results align with the averaged results presented in the main text.
@@ Appendix 0.D Per-scene Breakdown
Table 9: Comparison of point cloud quantities under different aggregation strategies on the Real Forward-facing dataset. For downsampling, we employ widely-used voxel downsampling, with a voxel size set to $2$. The iteration number for all strategies is set to $2$k.
@@
Figure 8: Depth maps visualization.
@@
Figure 10: Point cloud visualization.

### v3

@@ 2 Related Work
Generalizable NeRF. By implicitly representing scenes as continuous color and density fields using MLPs, Neural Radiance Fields (NeRF) achieve impressive rendering results with volume rendering techniques. Follow-up works [5, 46, 2, 30, 45, 38, 18, 39, 17] extend it to various tasks and achieve promising results. However, they all require time-consuming per-scene optimization. To address this issue, some generalizable NeRFs have been proposed. The general paradigm involves encoding features for each 3D point and then decoding these features to obtain volume density and radiance. According to the encoded features, generalizable NeRFs can be categorized into appearance features [55], aggregated multi-view features [40, 22, 24, 36], cost volume-based features [6, 26, 22, 24], and correspondence matching features [9]. Despite considerable progress, performance remains limited, with slow optimization and rendering speeds.
3D Gaussian Splatting (3D-GS) utilizes anisotropic Gaussians to explicitly represent scenes and achieves real-time rendering through differentiable rasterization. Motivated by this, several studies have applied it to various tasks, e.g. editing [8, 3], dynamic scenes [50, 27, 44], avatars [16, 15, 32] and others [7]. However, the essence of Gaussian Splatting still lies in overfitting the scene. To remedy this, a few concurrent works make initial attempts to generalize Gaussian Splatting to unseen scenes. The goal of Generalizable Gaussian Splatting is to predict Gaussian parameters in a feed-forward manner instead of per-scene optimization. PixelSplat [4] addresses scale ambiguity by leveraging an epipolar Transformer to encode features and subsequently decode them into Gaussian parameters. However, it focuses on image pairs as input and the Transformer incurs significant computational costs. GPS-Gaussian [57] draws inspiration from stereo matching and performs epipolar rectification and disparity estimation on input image pairs. However, it focuses on human novel view synthesis and requires ground-truth depth maps. Spatter Image [35] introduces a single-view 3D reconstruction approach. However, it focuses on object-centric reconstruction rather than generalizing to unseen scenes. Overall, these methods are constrained by inefficiency, limited to object reconstruction, and restricted to either image pairs or a single view. To this end, in this paper, we aim to study an efficient generalizable Gaussian Splatting for novel view synthesis in unseen general scenes.
@@ 3 Preliminary
$C=\sum_{i}c_{i}\alpha_{i}{\prod_{j=1}^{i-1}(1-\alpha_{i})}\,,$ (2)
where $c_{i}$ is the color of each point, defined by spherical harmonics (SH) coefficients. The density $\alpha_{i}$ is computed by the multiplication of 2D Gaussians and a learnable point-wise opacity. During optimization, the learnable attributes of each Gaussian are updated through gradient descent, including 1) a 3D position $\mu\in\mathbb{R}^{3}$, 2) a scaling vector $s\in\mathbb{R}_{+}^{3}$, 3) a quaternion rotation vector $r\in\mathbb{R}^{4}$, 4) a color defined by SH $c\in\mathbb{R}^{k}$ (where k is the freedom), and 5) an opacity $\alpha\in[0,1]$. Additionally, an adaptive density control module is introduced to improve rendering quality, comprising mainly the following three operations: 1) split into smaller Gaussians if the magnitude of the scaling exceeds a threshold, 2) clone if the magnitude of the scaling is smaller than a threshold, and 3) prune Gaussians with excessively small opacity or overly large scaling magnitudes.
@@ 4 MVSGaussian
Figure 2: Overview of MVSGaussian. We first extract features $\{f_{i}\}_{i=1}^{N}$ from input source views $\{I_{i}\}_{i=1}^{N}$ using FPN. These features are then aggregated into a cost volume, regularized by 3D CNNs to produce depth. Subsequently, for each 3D point at the estimated depth, we use a pooling network to aggregate warped source features, obtaining the aggregated feature $f_{v}$. This feature is then enhanced using a 2D UNet, yielding the enhanced feature $f_{g}$. $f_{g}$ is decoded into Gaussian parameters for splatting, while $f_{v}$ is decoded into volume density and radiance for depth-aware volume rendering. Finally, the two rendered images are averaged to produce the final rendered result.
@@ 4.3 Consistent Aggregation for Per-Scene Optimization
Figure 3: Consistent aggregation. With depth maps and point clouds produced by the generalizable model, we first conduct geometric consistency checks on depths to derive masks for filtering out unreliable points. The filtered point clouds are then concatenated to obtain a point cloud, serving as the initialization for per-scene optimization.
@@ 5.2 Generalization Results
5.3 Per-Scene Optimization Results
The quantitative results after per-scene optimization are reported in Table 3. For per-scene optimization, one strategy is to optimize the entire pipeline, similar to NeRF-based methods. Another approach is to optimize only the initial Gaussian point cloud provided by the generalizable model. When optimizing the entire pipeline, our method can achieve better performance with faster inference speeds compared to previous generalizable NeRF methods, and results comparable to NeRF, demonstrating the robust representation capabilities of our method. In contrast, optimizing only the Gaussians can significantly improve optimization and rendering speed because it eliminates the time-consuming feed-forward neural network. Moreover, performance can benefit from the adaptive density control module described in Sec. 3. Due to the excellent initialization provided by the generalizable model and the effective aggregation strategy, we achieve optimal performance within a short optimization period, approximately one-tenth of that of 3D-GS. Especially on the Real Forward-facing dataset, our method achieves superior performance with only $45$ seconds of optimization, compared to $10$ minutes for 3D-GS and $10$ hours for NeRF. Additionally, our method’s inference speed is comparable to that of 3D-GS and significantly outperforms NeRF-based methods. As shown in Fig. 5, our method is capable of producing high-fidelity views with finer details.
5.4 Ablations and Analysis
@@
Table 5: Comparison of different aggregation strategies. We report the quantitative results obtained with different strategies on the Real Forward-facing dataset [28]. For downsampling aggregation, we employ widely-used voxel downsampling with a voxel size set to $2$. The iteration number for all aggregation strategies is set to $2.5$k.
@@ 6 Conclusion
We present MVSGaussian, an efficient generalizable Gaussian Splatting approach. Specifically, we leverage MVS to infer depth, establishing a pixel-aligned Gaussian representation. To enhance generalization, we propose a hybrid rendering approach that integrates depth-aware volume rendering. Besides, thanks to high-quality initialization, our models can be fine-tuned quickly for specific scenes. Compared with generalizable NeRFs, which typically require minutes of fine-tuning and seconds of rendering per image, MVSGaussian achieves real-time rendering with superior synthesis quality. Moreover, compared with 3D-GS, MVSGaussian achieves better view synthesis with reduced training time.
@@ 21       $Mask\leftarrow Mask\cup Mask_{n}$
Implementation Details. Following ENeRF [22], we partition the DTU [1] dataset into $88$ training scenes and $16$ test scenes. We train the generalizable model on four RTX $3090$ GPUs using the Adam [20] optimizer, with an initial learning rate set to $45e-$ 4. The learning rate is halved every $50$k iterations. During the training process, we select $2$, $3$, and $4$ source views as inputs with respective probabilities of $0.1$, $0.8$, and $0.1$. For evaluation, we follow the criteria established in prior works such as ENeRF [22] and MVSNeRF [6]. Specifically, for the DTU test set, segmentation masks are employed to evaluate performance, defined based on the availability of ground-truth depth at each pixel. For Real Forward-facing dataset [28], where the marginal region of images is typically invisible to input images, we evaluate the $80\%$ area in the center of the images. This evaluation methodology is also applied to the Tanks and Temples dataset [21]. The image resolutions of the DTU, the Real Forward-facing, the NeRF Synthetic [29], and the Tanks and Temples datasets are $512\times 640$, $640\times 960$, $800\times 800$, and $640\times 960$ respectively. As discussed in Sec. 4.3 of the main text, we employ a consistency check to filter out noisy points for high-quality initialization. Specifically, we apply a dynamic consistency checking algorithm [47, 25], the details of which are provided in Algorithm 1. The predefined thresholds $\{\theta_{p}(n)\}_{n=1}^{N_{\theta}}$ are set to $\{\frac{n}{8}\}_{n=1}^{N_{\theta}}$, and $\{\theta_{d}(n)\}_{n=1}^{N_{\theta}}$ are set to $\{\frac{n}{10}\}_{n=1}^{N_{\theta}}$. For 3D-GS [19], following [58], we use COLMAP [33] to reconstruct the point cloud from the working set (training views) as initialization. Specifically, we employ COLMAP’s automatic reconstruction to achieve the reconstruction of sparse point clouds. Some examples are shown in Fig. 7. As mentioned in Sec.5.1 of the main text, our optimization strategy and hyperparameters settings remain consistent with the vanilla 3D-GS, except for the number of iterations. The iterations of our method on Real Forward-facing, NeRF Synthetic and Tanks and Temples datasets are $2.5$k, $5$k and $5$k, respectively.
@@
Appendix 0.B Additional Ablation Experiments
@@
Density for Volume Rendering. Since only one point per ray is sampled, our model predicts single radiance $r$ and density $\sigma$. In this case, the pixel’s color $c$ obtained through volume rendering is given by $c=(1-\exp(-\sigma))r$. This resembles pixel-aligned splatting, where one Gaussian contributes one pixel, sharing the alpha-based rendering principles but offering a simpler implementation. Therefore, predicting density is necessary as it indicates the point’s opacity, as validated by ablation results in Table 8.
Table 8: The ablation study on the density prediction.
Settings DTU Real Forward-facing NeRF Synthetic
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
w/o density 28.03 0.963 0.076 23.96 0.854 0.165 26.22 0.947 0.071
w density 28.21 0.963 0.076 24.07 0.857 0.164 26.46 0.948 0.071
Initialization Comparison. Our generalization model can provide a denser point cloud for 3D-GS [19] than Structure-from-Motion (SfM) as initialization. Considering that the MVS method can also obtain a denser point cloud, we conduct the comparison in Table 9. MVS methods are typically trained on DTU [1] and BlendedMVS [53], then tested on Tanks and Temples dataset [21]. Thus, we select the latest ET-MVSNet [25] and compare it on Tanks and Temples dataset. While ET-MVSNet [25] surpasses SfM, it’s still limited. Because it focuses solely on accurate depth, while our method generates point clouds tailored for view synthesis. Depth and view quality aren’t directly proportional, as mentioned in previous works such as ENeRF [22].
Table 9: Initialization Comparison.
Initialization PSNR SSIM LPIPS Timeft FPS
ET-MVSNet [25] 22.66 0.861 0.204 90s 300$+$
Ours 24.58 0.903 0.137 90s 300$+$
@@
Table 11: Comparison of point cloud quantities under different aggregation strategies on the Real Forward-facing dataset. For downsampling, we employ widely-used voxel downsampling, with a voxel size set to $2$. The iteration number for all strategies is set to $2.5$k.
@@
Inference Speed Analysis. As shown in Table 1 of the main text, the inference speed (FPS) of our generalizable model is $21.5$. Here, we present the inference time breakdown result in Table 12. The primary time overhead comes from the neural network, while the subsequent rendering process incurs minimal time overhead. Therefore, we discard the neural network component during the per-scene optimization stage, resulting in a significant increase in speed.
Appendix 0.C More Qualitative Results
Appendix 0.D Per-scene Breakdown
As shown in Tables 131615, and 14, we present the per-scene breakdown results of DTU [1], NeRF Synthetic [29], Real Forward-facing [28], and Tanks and Temples [21] datasets. These results align with the averaged results presented in the main text.
Figure 8: Depth maps visualization. We visualize the depth maps predicted by our method on different datasets [1, 29, 28].
@@ Appendix 0.D Per-scene Breakdown
Figure 10: Point cloud visualization under different aggregation strategies.